Given the following:
C = 2000 + 0.75Yd
T = 300
I = 320
G = 300
X = 300
M=100


a) Determine the equilibrium level of income using expenditure and injection-leakage approach

b) Determine the value of C at equilibrium level of income.

c) Calculate the equilibrium level of income when there is an increase in investment of 100 using expenditure and multiplier approach.
1
Expert's answer
2020-06-21T19:54:07-0400

a) Determine the equilibrium level of income using expenditure and injection-leakage approach


We have the data:


"C = 2000 + 0.75Yd\\[0.3cm]

T = 300\\[0.3cm]

I = 320\\[0.3cm]

G = 300\\[0.3cm]

X = 300\\[0.3cm]

M=100"


The expenditure approach method is:



"Y= C + I + G + (X - M)\\[0.3cm]"

Therefore:



"Y = 2000 + 0.75(Y - 300) + 320 + 300 + (300 - 100)\\[0.3cm]

Y = 2820+ 0.75Y - 225 \\[0.3cm]
0.25Y = 2595\\[0.3cm]
Y^* = \dfrac{2595}{0.25} = \color{blue}{10,380}"

b) Determine the value of C at equilibrium level of income.



"C = 2000 + 0.75(10,380 - 300)\\[0.3cm]
\color{red}{C = 9,560}"

c) Calculate the equilibrium level of income when there is an increase in investment of 100 using expenditure and multiplier approach.


The multiplier is given by:



"k = \dfrac{\Delta Y}{\Delta I} = \dfrac{1}{1 - MPC}"

In our question, MPC = 0.75. Therefore:



"k = \dfrac{1}{1 - 0.75} = 4"

When the investment increases by 100, the equilibrium income will increase by:



"\Delta Y = 4\times \Delta 100\\[0.3cm]
\Delta Y = 400"

The new income is:



"Y^{**} = 10,380 + 400 = \color{blue}{10,780}"